OpenStudy (anonymous):

whats the vertex of f(x)=x^2-2x-8

OpenStudy (anonymous):

change it to the completed square form... can u?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first coordinate of the vertex is alway \(-\frac{b}{2a}\) in your case it is \(-\frac{-2}{2\times 1}=1\) the second coordinate of the vertex is what you get when you replace \(x\) by that number

OpenStudy (raden):

1^2-2(1)-8=...

OpenStudy (anonymous):

the completed square form is a(x-h)^2+k

OpenStudy (anonymous):

try to take it to that form